OpenStudy (anonymous):

HELP PLEASE Juanelda's doctor has prescribed medication that she must take every 5 hours. She started the medication at 10:00 a.m. on Thursday. What is the first day on which she will take her medication at noon? Write an equation that will provide this answer.

OpenStudy (***[isuru]***):

this is hard ...

OpenStudy (***[isuru]***):

I got it bro

OpenStudy (***[isuru]***):

Ok , let's name of the person as "A" cause her name is troublesome The question asks when will "A" take medicine at noon. "A" takes her first dose at 10.a.m. so she can't take the medicine at noon on Thursday cause there r only 2 hours left for noon

OpenStudy (***[isuru]***):

So the day she take the medicine at noon is not thursday

OpenStudy (***[isuru]***):

It has to be a day which comes after Thursday. Also If we start counting from Thursday noon the nearest possible day which satisfy the question is Friday noon

OpenStudy (***[isuru]***):

Which mean exactly 24 hours after Thursday noon

OpenStudy (***[isuru]***):

If "A" starts take medicine on Thursday noon the total time until he will spend until he take a medicine at noon will be a multiply of 24 So if n = the total no.of medicine "A" has take 5*n/24 = N (a whole number )

OpenStudy (***[isuru]***):

but that will happen only if "A" takes her pill at Thursday noon. But here she take the 1st pill at 10 a.m so u have to subtract 2 hours in order to get a multiply of 24

OpenStudy (***[isuru]***):

So the final answer is like this \[\frac{ 5n - 2 }{ 24} = \mathbb{N} \] the first value which will satisfy the above expression is the key to the answer here , 10 is the value for n So "A" will take his 10 th pill at noon the day is saturday

OpenStudy (anonymous):

Thank you so much

OpenStudy (***[isuru]***):

u r welcome bro

OpenStudy (anonymous):

that was a nice question bro...!!